OpenStudy (anonymous):

If sin theta< 0 and tan theta< 0 then: A. 0 < theta < 90 B. 90 < theta < 180 C. 180 < theta < 270 D. 270 < theta < 360

OpenStudy (jdoe0001):

\(\bf sin theta< 0 and tan theta< 0 then:\\ sin(\theta)<0\qquad \textit{another way of saying }sin(\theta)\textit{ is negative}\\ \quad \\ tan(\theta) <0 \textit{another way of saying }tan(\theta)\textit{ is negative}\\ \quad \\ tan(\theta)=\cfrac{sin(\theta)}{cos(\theta)}\implies \cfrac{{\color{red}{ negative}}}{{\color{blue}{ positive}}}\implies {\color{blue}{ cos(\theta) > 0}}\textit{ or is positive}\) now... check your Unit Circle to see where the cosine and sine are such values

OpenStudy (jdoe0001):

woops I pasted too much but anyhow \(\bf sin(\theta)<0\qquad \textit{another way of saying }sin(\theta)\textit{ is negative}\\ \quad \\ tan(\theta) <0 \textit{another way of saying }tan(\theta)\textit{ is negative}\\ \quad \\ tan(\theta)=\cfrac{sin(\theta)}{cos(\theta)}\implies \cfrac{{\color{red}{ negative}}}{{\color{blue}{ positive}}}\implies {\color{blue}{ cos(\theta) > 0}}\textit{ or is positive}\) check your Unit Circle

OpenStudy (cwrw238):

Hint both sine and tangent are negative in the fourth quadrant

OpenStudy (anonymous):